Calculate the molarity of 0.060 moles of NaHCO3 in 1500. mL of solution
1
Expert's answer
2020-05-07T14:15:12-0400

C(NaHCO3) = n(NaHCO3)/V(NaHCO3) = 0.060/1.5 = 0.04 M